OpenStudy (anonymous):

Find the identical expression for the following. 1-sin^4x\1+sin^2x

hartnn (hartnn):

write 1 = 1^2 and \(\sin^4x=(\sin^2x)^2\) then, use the formula \(a^2-b^2=(a+b)(a-b)\) then see what gets cancelled!

OpenStudy (campbell_st):

just factor the numerator, its the difference of 2 squares

OpenStudy (anonymous):

thank you